OpenStudy (anonymous):

2sqrt50 - 3sqrt125 +sqrt98

OpenStudy (campbell_st):

\[2\sqrt{50} = 2\sqrt{25\times2} = 10\sqrt{2}\] \[3\sqrt{125} = 3\sqrt{25\times5} = 15\sqrt{5}\] \[\sqrt{98} = \sqrt{49\times2} = 7\sqrt{2}\] now collect the like terms tp get the answer